OpenStudy (anonymous):

Choose the correct simplification of the expression (3x2y3z4)(2x5y2z3). 6x7y5z7 6x10y6z12 5 x7y5z7 5x10y6z12

OpenStudy (anonymous):

If I'm not mistaking is 6x10y6z12

OpenStudy (anonymous):

thanks for trying <3

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

6x7y5z7 b/c you add the exponent degree when multiplying x(x) = x^1+1 = x^2 (x^2)^5= (x)^5*2 =x^10

OpenStudy (anonymous):

idk